OpenStudy (anonymous):

f(x)=x^3+6x^2+9x find: x intercepts vertical asymptotes horizontal/slant asymptotes derivative second derivative increasing interval decreasing interval concave up interval concave down interval

OpenStudy (dumbcow):

=x(x^2 +6x+9) = x(x+3)^2 = 0 intercepts: x=0,-3 there are no asymptotes since it is a polynomial f'(x) = 3x^2 +12x +9 ....(using power rule x^n -> n*x^(n-1) ) f''(x) = 6x +12 function is increasing when f'(x) > 0 function is decreasing when f'(x) < 0 function is concave up when f''(x) > 0 function is concave down when f''(x) < 0

OpenStudy (anonymous):

thank you!